\begin{thebibliography}{AAAA}
%\bibitem[Bl]{Beilinson} A. Beilinson,
%\emph{Remarks on Grothendieck's standard conjectures}, Regulators, 25--32,
%Contemp. Math., 571, Amer. Math. Soc., Providence, RI, 2012.

\bibitem[AC]{Akyildiz-Carrell}
E. Akyildiz and J. Carrell,
\emph{A generalization of the Kostant-Macdonald identity}.
Proc. Nat. Acad. Sci. U.S.A. 86 (1989), no. 11, 3934--3937.

\bibitem[Be]{Beauville} A.Beauville,
\emph{Compact complex surfaces}, 2nd ed., London Mathematical Society Student Texts, vol. 34, Cambridge University Press, Cambridge, England, 1996. Translated from the 1978 French original by R. Barlow, with assistance from N. I. Shepherd-Barron and M. Reid.

\bibitem[B-B1]{Bialynicki-Birula1}
A. Bia{\l}ynicki-Birula,
{\sl On fixed point schemes of actions of multiplicative and additive groups.}
Topology 12 (1973), 99--103.

\bibitem[B-B2]{Bialynicki-Birula2}
A. Bia{\l}ynicki-Birula,
{\sl Some theorems on actions of algebraic groups.}
Ann. of Math. (2) 98 (1973), 480--497.


\bibitem[C]{Carrell}
J. Carrell,
\emph{Holomorphic $\C^*$ actions and vector fields on projective varieties}. Topics in the theory of algebraic groups, 1--37,
Notre Dame Math. Lectures, 10, Univ. Notre Dame Press, Notre Dame, IN, 1982.

\bibitem[CG]{Carrell-Goresky} J. B. Carrell and R. M. Goresky,
{\sl A decomposition theorem for the integral homology of a variety.}
Invent. Math. 73 (1983), no. 3, 367--381.

\bibitem[CHK]{Carrell-Howard-Kosniowski}
J.  Carrell, A. Howard and C. Kosniowski,
{\sl Holomorphic vector fields on complex surfaces.}
Math. Ann. 204 (1974), 73--82.

\bibitem[CL]{Carrell-Lieberman}
J. Carrell and D. Lieberman,
Holomorphic vector fields and Kaehler manifolds.
Invent. Math. 21 (1973), 303--309.



\bibitem[CS1]{Carrell-Sommese1}J. B. Carrell and A. J. Sommese,
{\sl ${\C}^{\ast}$-actions. } Math. Scand. 43 (1978/79), no. 1, 49--59.

\bibitem[CS2]{Carrell-Sommese2} J. B. Carrell and A. J. Sommese,
{\sl Some topological aspects of $\C^{\ast} $ actions on compact Kaehler manifolds.}
Comment. Math. Helv. 54 (1979), no. 4, 567--582.

\bibitem[ES]{Ellingsrud-Stromme} G. Ellingsrud and S. A. Stromme,\emph{ Toward the Chow ring of the Hilbert scheme of $\P^2$},
J. Reine Agnew. Math. 441 (1993), 33--44.


%\bibitem[GKZ]{Gelfand-Kapranov-Zelevinsky}
%I. M. Gel'fand, M. M. Kapranov and A.V. Zelevinsky,
%\emph{Discriminants, resultants, and multidimensional determinants}.
%Mathematics: Theory \& Applications. Birkh\"{a}user Boston, Inc., Boston, MA, 1994.


\bibitem[Ha]{Hartshorne}
R. Hartshorne,
\emph{Algebraic geometry}.
Graduate Texts in Mathematics, No. 52. Springer-Verlag, New York-Heidelberg, 1977.


\bibitem[HL]{Harvey-Lawson}
F. R. Harvey and H. B. Lawson, Jr.
{\sl Finite volume flows and Morse theory.}
Ann. of Math. (2) 153 (2001), no. 1, 1--25.


%\bibitem[Ho]{Horrocks}
%G. Horrocks,
%{\sl Fixed point schemes of additive group actions.}
%Topology 8 1969 233--242.

%\bibitem[Hu1]{Hu} W. Hu, {\sl Homotopy groups of Chow varieties.}  Amer. J. Math.,137(2015), no.6, 1431--1440.

%\bibitem[Hu2]{Hu2} W. Hu,
%{\sl Algebraic cycles of a fixed degree.} Proc. Amer. Math. Soc.,138 (2010), 2365--2373.

%\bibitem[Hu3]{Hu3}W. Hu,
%\emph{The multiplicative group action on singular varieties and Chow varieties}. Preprint.

\bibitem[Hw1]{Hwang1}
J. M. Hwang,
{\sl Additive vector fields, algebraicity and rationality}.
Math. Ann. 304 (1996), no. 4, 757–767.

\bibitem[Hw2]{Hwang2}
J. M. Hwang,
{\sl Holomorphic vector fields with totally degenerate zeroes}.
Compositio Math. 105 (1997), no. 1, 65–77.

\bibitem[Fr]{Frankel}
T. Frankel,
{\sl  Fixed points and torsion on K\"ahler manifolds}
Ann. of Math. (2) 70 1959 1--8.

%\bibitem[FL]{Friedlander-Lawson}
%E. Friedlander and  H. B. Lawson, Jr.
%\emph{A theory of algebraic cocycles}.
%Ann. of Math. (2) 136 (1992),  no. 2, 361--428.

\bibitem[La]{Lawson1} H. B. Lawson, Jr.
{\sl Algebraic cycles and homotopy
theory.} Ann. of Math. {\bf 129}(1989), 253-291.

%\bibitem[Le]{Lehmann}
%B. Lehmann,
%{\sl Asymptotic behavior of the dimension of the Chow variety.}
%Adv. Math. 308 (2017), 815--835.


\bibitem[Li1]{Lieberman1}
D. I. Lieberman,
{\sl Holomorphic vector fields on projective varieties.} Several complex variables (Proc. Sympos. Pure Math., Vol. XXX, Part 1,
Williams Coll., Williamstown, Mass., 1975), pp. 273--276. Amer. Math. Soc., Providence, R.I., 1977.

\bibitem[Li2]{Lieberman2}
D. I. Lieberman,
{\sl Holomorphic vector fields and rationality. Group actions and vector fields} (Vancouver, B.C., 1981), 99--117,
Lecture Notes in Math., 956, Springer, Berlin, 1982.



\bibitem[LF]{Lima-Filho} P. Lima-Filho, {\sl Lawson homology for
quasiprojective varieties.}  Compositio Math. 84(1992), no. 1,
1--23.

%\bibitem[M]{McDuff}
%D. McDuff, {\sl  The moment map for circle actions on symplectic manifolds}.
%J. Geom. Phys. 5 (1988), no. 2, 149--160.

\bibitem[K]{Konarski}
J. Konarski,
{\sl On a special case of Carrell's conjecture}.
J. Algebra 182 (1996), no. 1, 38--44.



\bibitem[PS]{Parshin-Shafarevich}
{\sl Algebraic geometry.} V.
Fano varieties. Edited by A. N. Parshin and I. R. Shafarevich. Encyclopaedia of Mathematical Sciences, 47. Springer-Verlag, Berlin, 1999.


%\bibitem[RY]{Reichstein-Youssin}
%Z. Reichstein and B. Youssin,
%{\sl Equivariant resolution of points of indeterminacy}. Proc. Amer. Math. Soc. 130 (2002), no. 8, 2183--2187.

%\bibitem[Sh]{Shafarevich}
% I. R. Shafarevich,
%\emph{Basic algebraic geometry.}
%Springer Study Edition. Springer-Verlag, Berlin-New York, 1977.

\bibitem[Si]{Siu}
Y.T. Siu,
{\sl Global nondeformability of the complex projective space.} Prospects in complex geometry (Katata and Kyoto, 1989), 254--280,
Lecture Notes in Math., 1468, Springer, Berlin, 1991.

\bibitem[So]{Sommese}
A. J. Sommese,
\emph{Extension theorems for reductive group actions on compact Kaehler manifolds}.
Math. Ann. 218 (1975), no. 2, 107--116.

%\bibitem[Su]{Sumihiro}
%H. Sumihiro,
%{\sl Equivariant completion}.  J. Math. Kyoto Univ., 14, no. 1 (1974), 1--28.

%\bibitem[V]{Voisin}
%C. Voisin,
%\emph{Hodge theory and complex algebraic geometry.} II.
%Translated from the French by Leila Schneps. Cambridge Studies in Advanced Mathematics, 77. Cambridge University Press, Cambridge, 2007.

\end{thebibliography}

\documentclass[9pt]{amsart}

\usepackage{amssymb,latexsym,amsmath}
\usepackage{graphics}                 
\usepackage{color}                    
\usepackage{hyperref}                 
\usepackage[all]{xy}

\numberwithin{equation}{section}
\theoremstyle{plain} 
\newtheorem{theorem}[equation]{Theorem}
\newtheorem{corollary}[equation]{Corollary}
\newtheorem{lemma}[equation]{Lemma}
\newtheorem{proposition}[equation]{Proposition}

\theoremstyle{definition}

\newtheorem{example}[equation]{Example}
\newtheorem{conjecture}[equation]{Conjecture}

\theoremstyle{remark}
\newtheorem{remark}[equation]{Remark}

\def\C{{\mathbb{C}}}

\def\P{{\mathbb{P}}}

\title{Holomorphic vector fields and rationality}

\author{Wenchuan Hu}
\keywords{Holomorphic vector field,  rationality}

\address{
School of Mathematics\\
Sichuan University\\
Chengdu 610064\\
P. R. China
}

\email{huwenchuan@gmail.com}

\begin{document}

\begin{abstract}
We show that a nonsingular  complex projective  variety  admitting a holomorphic
vector field with nonempty isolated zeroes, is rational  using a key  technique by Harvey-Lawson on finite volume flows.
This statement was conjectured by J. Carrell. By the same technique, we obtain a uniform  upper bound of Betti numbers of
nonsingular complex projective  variety  admitting a holomorphic
vector field with exact one zero point. Such an upper bound  depends only on the dimension of the variety, which is a stronger
version of a result of Akyildiz and Carrell.

\end{abstract}

\maketitle
\pagestyle{myheadings}
 \markright{Holomorphic vector fields and rationality}


\section{Introduction}
The action of  group actions on manifolds is one of long lasting research subject in mathematics.
Almost every  topic in algebra, geometry and topology, there is a corresponding version with group actions, e.g., for a cohomology theory
there is an equivariant cohomology theory. In particular, the action of $1$-parameter group
on  manifolds is one of the simplest and widest concerned interesting branches in the field of group actions on topological spaces.



There are lots of significant results related  the structure of a complex analytic space or a complex manifold to its fixed point set
under the action of a complex $1$-parameter group.
The topological and analytic invariants  between a compact K\"{a}hler manifold or a singular projective variety
 and  those of  the fixed point set under an analytic action of
 a complex multiplicative group  can be found in \cite{Bialynicki-Birula1},\cite{Carrell-Goresky},\cite{Frankel},  etc. and the references therein.

Bialynicki-Birula showed that any nonsingular projective variety $X$ under a complex multiplicative group $G_m$
action admits a canonical decomposition into $G_m$-invariant locally closed subvarieties. In particular,
the algebro-geometric invariants of the variety can be read from those of the fixed point set. When the fixed point set is isolated,
all the cohomology theories on $X$ (e.g., Chow groups, algebraic cycles modulo algebraic
equivalence, Lawson homology, singular homology) are naturally isomorphic (see  \cite{Bialynicki-Birula1}, \cite{Ellingsrud-Stromme}, \cite{Frankel}, \cite{Lima-Filho}).
Moreover, such a variety $X$ must be a rational variety.

However, the situation is much more subtle for a projective variety under the action of an algebraic  additive group.
The fact is that the understanding of the structure of a nonsingular complex projective variety $X$ through
the fixed point set under an additive group action  amounts to the understanding of the structure of $X$ through the zero locus of a
holomorphic vector field. The reason is that the structure of a nonsingular projective variety under
the action of multiplicative group is clearly
understood through the information of the fixed point set (see \cite{Lieberman2}).

Now let $V$ be a holomorphic vector field defined on a nonsingular complex projective algebraic variety
$X$. The zero subscheme  $Z$ is the subspace of $X$ defined by the ideal generated by $V\mathcal{O}_X$.

 Contrary to the real case,
 there are a lot of topological obstructions  to the existence of a holomorphic vector field on
 a complex projective variety.  Moreover, the structure of $X$  are closely related to that of $Z$ (see \cite[\S 5]{Carrell}).
  For examples, the rational cohomological ring  of $X$ can be determined by those of $Z$
 and the fundamental group of $X$ is perfect if $Z$ is simple (i.e., smooth
of dimension zero) (see \cite{Carrell-Lieberman}).


 A fundamental problem in algebraic geometry  is to determine which varieties are rational.
One of the most famous problems in the direction, formulated by J. Carrell,  is the following conjecture.

\begin{conjecture}[\cite{Carrell}, p.33]\label{Conj1}
A smooth (or nonsingular) irreducible  projective algebraic variety $X$ admitting a holomorphic
vector field with nonempty isolated zeroes, is rational.
\end{conjecture}

This conjecture  stated in an open problem format, can be found in an earlier literature (see  \cite{Lieberman1}).

During last fourty years, there have been lots of progress in all kinds of special cases.
For surfaces, this conjecture is proved in \cite{Carrell-Howard-Kosniowski}, \cite[p.99]{Lieberman2}.
For $X$ with $\dim X\leq 5$, it was proved by Hwang (see \cite{Hwang1}).
 Hwang also mentioned in his paper that M. Koras had proved the conjecture  for
dimension less than or equals to 4 in an unpublished paper in a possibly quite different way.
Under the assumption that the automorphism group of the manifold is semisimple, it was proved by Deligne (see \cite[p.32]{Carrell}).
Under the assumption that the induced action on the tangent space at the isolated zero has a single
Jordan block, it was proved by  Konarski ([K]). Under the assumption that  the holomorphic vector field  has a totally degenerate zero,
it was proved by  Hwang (\cite{Hwang2}).
This conjecture has been studied by Lieberman (see \cite{Lieberman1}, \cite{Lieberman2}),
who had shown that it holds if every point in $Z$ is
non-degenerate. Recall that a  point in $Z$ is called non-degenerate if
it is a nonsingular point on $Z$ and  the $\mathcal{O}_{X, Z}$-linear map $L_V:I_Z/I_Z^2\to I_Z/I_Z^2$ has nonzero determinant on $Z$.
Note that a holomorphic vector field obtained from the flow of a multiplicative group action is always non-degenerate.

A stronger version of the conjecture (might be also made by Carrell) that every smooth projective variety
admitting a holomorphic vector field with nonempty
isolated zeros also admits a $\C^*$-action with
nonempty isolated zeros.

Conjecture \ref{Conj1}, as is well known,  is equivalent to the following(\cite{Lieberman1},\cite{Konarski}):
\begin{conjecture}\label{Conj2}
If $X$ is a nonsingular projective variety admitting an algebraic additive group
action having exactly one fixed point, then $X$ is rational.
\end{conjecture}

We will show below this conjecture to be true by applying techniques developed by Harvey and Lawson to the ring  $\mathcal{O}_{X,p}$ of local regular functions of a point on $X$, to get an isomorphism
between $\mathcal{O}_{X,p}$ and that of a point on $\C^n$.

As  applications of  techniques developed by Harvey and Lawson, we obtain a uniform upper bound $C_n=2^n$ for the Euler number
 and Hodge numbers of $X$ of dimension $n$, where $X$ is a nonsingular projective variety  admitting an algebraic additive group action with exactly one fixed point (see Theorem \ref{Th3.2}).

 In particular, we have  $b^2(X)\leq \dim X$ for such an $X$.
 This is a generalization of an ``unexpected consequence" of  Akyildiz and Carrell(see \cite[Th.2]{Akyildiz-Carrell}). In Akyildiz and Carrell theorem,  much more stronger conditions for $X$ are assumed.
  We give a classification of surfaces  and a partial  classification of threefolds  of  admitting a $\C$-action with exact one fixed point.

\section{Proof of Carrell's conjecture}
In this section we will show  Conjecture \ref{Conj2} and hence  Conjecture \ref{Conj1} hold.
From now on, let $X$ be a smooth complex projective variety of dimension $n$ admitting a $\C$-action with exactly one fixed point $p$, i.e, $X^{\C}=p$.

Let $m_p\subset \mathcal{O}_{X,p}$ (resp. $m_o\subset \mathcal{O}_{\C^n,o}$) be the corresponding maximal ideal  of local algebras.

Note that the  flow $\phi_t$ induced from the $\C$-action generates a finite volume complex graph $\mathcal{T}$ in the sense of Harvey and Lawson (see \cite[\S9]{Harvey-Lawson}), where
$$
\mathcal{T}:=\{(t,\phi_t(x),x)\in \C\times X\times X|t\in \C  ~{\rm and}~ x\in X \}\subset \P^1\times X
\times X
$$
 and its closure $\overline{\mathcal{T}}\subset  \P^1\times X \times X$ is a projective variety.  According to their theory, this $\overline{\mathcal{T}}$ provides us an operator equation
 \begin{equation}\label{eq2.1}
 I-P=d\circ T+T\circ d:\mathcal{E}^k(X)\to \mathcal{D}'^k(X)
 \end{equation}
for $k\geq 0$, where $\mathcal{E}^k(X)$ is the space of  smooth $k$-forms on $X$, $\mathcal{D}'^k(X)$ is the space of $k$-currents on $X$,
 $I$ is the identity operator, $P:\mathcal{E}^k(X)\to \mathcal{D}'^k(X)$ is given by $P\alpha=\lim_{t\to \infty}\phi_t^*(\alpha)$ and $T:\mathcal{E}^k(X)\to \mathcal{D}^{k-1}(X)$
is the operator given by $\overline{\mathcal{T}}$ (see \cite[\S 2]{Harvey-Lawson}).

When restricting to the local ring of $\C$-algebra $\mathcal{O}_{X,p}$, we get the formula
\begin{equation}\label{eqn2.2}
 I-P=T\circ d:\mathcal{O}_{X,p}\to \mathcal{O}_{X,p}
 \end{equation}
since $T=0$ on $\mathcal{E}^0(X)=C^{\infty}(X)$ the set of smooth functions on $X$ and hence $T=0$ on $\mathcal{O}_{X,p}$
since an element $f$ in $\mathcal{O}_{X,p}$ can be extended to a smooth function on $X$.
The operator $P$ on $f\in C^{\infty}(X)$ is the constant function with the value $f(p)$  since
 $$(Pf)(x)=\lim_{t\to \infty} \phi_t^*(f)(x)=\lim_{t\to \infty}f\circ \phi_t(x)=f(\lim_{t\to \infty}\phi_t(x))=f(p).$$

The details are given in the following lemma, where Equation \eqref{eqn2.2} is applied to  local functions in $\mathcal{O}_{X,p}$.
\begin{lemma}\label{lemma2.2}
One gets the following formula
\begin{equation}\label{eqn2.4}
f-f(p)=T\circ df, f\in \mathcal{O}_{X,p}.
\end{equation}
\end{lemma}
\begin{proof}
Let $F,\tilde{F}\in C^{\infty}(X)$ be two extensions of the germ $f\in \mathcal{O}_{X,p}$. Then
we have $F-F(p)=T\circ dF$ and $\tilde{F}-\tilde{F}(p)=T\circ d\tilde{F}$ by Equation \eqref{eq2.1}. Since $F(p)=\tilde{F}(p)=f(p)$, we get
$F-\tilde{F}=T\circ d(F-\tilde{F})$. Since both $F$ and $\tilde{F}$ are the extension of $f$, there exists a neighborhood $U_p$ of $p$ such that
$F-\tilde{F}\equiv 0$ on $U_p$.  That is, the operator $T\circ d$ applies to $F-\tilde{F}$ is zero on on $U_p$.
Hence two different choices of extensions of $f\in \mathcal{O}_{X,p}$ yield the same local function whose
class is $f-f(p)$. This shows that $I-P=T\cdot d$ holds on $ \mathcal{O}_{X,p}$.
\end{proof}

Moreover, when restricted on $m_p$, we have
\begin{equation}\label{eqn2.5}
f=T\circ df, f\in m_p
\end{equation}
since  $f(p)=0$ for $f\in m_p$.

Let $\Omega^1_{X,p}$ be the space of local 1-forms with coefficients in $\mathcal{O}_{X,p}$, i.e., $\Omega^1_{X,p}=\{df|f\in \mathcal{O}_{X,p}\}$.
Since $X$ is a nonsingular variety, we have $T_p^*X\cong\Omega^1_{X,p}\otimes_{\mathcal{O}_{X,p}} \C\cong m_p/m_p^2$ as $\C$-vector spaces, where $T_p^*X$ is
the cotangent space of $X$ at $p$.

Since $p\in X$ is nonsingular, we have $m_p/m_p^2\cong T^*_pX$ as $\C$-vector spaces. Since $\dim_{\C} X=n$, there exists a $\C$-linear isomorphism
 $\psi^*:T^*_pX \stackrel{\cong}{\to} \C^n$. Let $\omega_i$($1\leq i\leq n$) be a basis of $T^*_pX$ such that $\psi^*(\omega_i)=dz_i$, where
 $z_i$($1\leq i\leq n$) is  a system of  local coordinates of $\C^n$ at the origin.

Set $y_i:=T(\omega_i)$, then $y_i\in m_p$ by Equation \eqref{eqn2.4}. Since $ T(dy_i)=y_i$  by Equation \eqref{eqn2.5}, we have $dy_i-\omega_i\in \ker(T)$.
We can choose $dy_1,...,dy_n$ as the basis of $ T^*_pX$. That is, we give a system of local coordinates $y_1,...,y_n$ for $X$ at $p$ such that
there exists a map $\psi: U_o\to U_p$ is defined  from a neighborhood $U_o \subset\C^n$ of $o$ to a neighborhood  $U_p\subset \in X$ of $p$.

 For $\omega\in \Omega^1_{X,p}$, we have $\omega=\sum_if_idy_i=df$ for some $f\in \mathcal{O}_{X,p}$, $T(\omega)=T(df)=f$, where
 $f_i=\frac{\partial f}{\partial y_i}$.

Now we extend the map $\psi^*:T^*_pX \stackrel{\cong}{\to} \C^n$ to $\psi^*:\Omega^1_{X,p}\to \Omega^1_{\C^n,o}$ as the following:
$$\psi^*(df)(z_1,...,z_n):=df(y_1,...,y_n), \forall df\in \Omega^1_{X,p}. $$

Now  we define a map $\psi^*:\mathcal{O}_{X,p}\to \mathcal{O}_{\C^n,o}$ (by the same notation) as  $\psi^*(f)(z_1,...,z_n):=f(y_1,...,y_n)$ for $f\in \mathcal{O}_{X,p}$.
This map $\psi^*:\mathcal{O}_{X,p}\to \mathcal{O}_{\C^n,o}$ is a local ring homomorphism since it is a $\C$-linear map and
$\psi^*(f_1f_2)=\psi^*(f_1)\psi^*(f_2)$. To see the last formula, we get from the definition of $\psi^*$ that
$\psi^*(f_1\cdot f_2)(z_1,...,z_n)= (f_1\cdot f_2)(y_1,...,y_n)=f_1(y_1,...,y_n)\cdot f_2(y_1,...,y_n)=\psi^*(f_1)(z_1,...,z_n)\cdot\psi^*(f_2)(z_1,...,z_n)$.

\begin{lemma}
This map $\psi^*:\mathcal{O}_{X,p}\to \mathcal{O}_{\C^n,o}$ defined by $\psi^*(f)(z_1,...,z_n)=f(y_1,...,y_n)$ for $f\in \mathcal{O}_{X,p}$ is injective.
\end{lemma}
\begin{proof}
Let $f\in \mathcal{O}_{X,p}$ such that $\psi^*(f)=0$. Then  $0=\psi^*(f)(z_1,...,z_n)=f(y_1,...,y_n)$ for $(y_1,...,y_n)$ in the neighborhood of $p$. This implies
that $f$ is zero in $\mathcal{O}_{X,p}$.
\end{proof}

Now we define map $T_o:\Omega^1_{\C^n,o}\to m_o$, $T_o(dg):=g$ for $g\in \mathcal{O}_{\C^n,o}$. Then $T_o(dz_i)=z_i$ for $1\leq i\leq n$.

\begin{lemma}\label{lemma2.6}
There exists a commutative diagram
\begin{equation}\label{eq2.7}
\xymatrix{
& {\mathcal O}_{X,p}\ar[r]^d\ar[d]^{\psi^*} & {\Omega}^1_{X,p}\ar[d]^{\psi^*}\ar[r]^{T}\ar[d]& {\mathcal O}_{X,p}\ar[d]^{\psi^*}\\
&{\mathcal O}_{\C^n,o}\ar[r]^d &  {\Omega}^1_{\C^n,o}\ar[r]^{T_o}& {\mathcal O}_{\C^n,o},
}
\end{equation}
Moreover, $\psi^*(m_p)\subset m_o$.
\end{lemma}

\begin{proof}
The first square in Equation \ref{eq2.7} is commutative since $d\circ \psi^*(f)= \psi^* \circ df$ by their definitions,
while  the second square in Equation \ref{eq2.7} is commutative since $T\circ d=Id-P$ and the definition of $T_o$.
To see this,  we take $\alpha=df$ where $f(p)=0$, then
$T_o(\alpha)=f$. Hence $\psi^*T(\alpha)=\psi^*(f)=T_o(d(\psi^*(f)))=T_o(\psi^*(df))=T_o(\psi^*(\alpha))$
 since the map $\psi^*: \mathcal{O}_{X,p}\to \mathcal{O}_{\C^n,o}$ and $\psi^*:\Omega^1_{X,p} \to \Omega^1_{\C^n,0}$ are natural maps.
\end{proof}

\begin{remark}
The key point in Lemma \ref{lemma2.6} is that there exists map $T_o:\Omega^1_{\C^n,o}\to m_o$ such that $T_o\circ d=I$ on $m_o$.
\end{remark}

Note that there are   natural $\C$-linear isomorphisms
\begin{equation}\label{eqn2.9}
\Omega^1_{X,p}\cong \bigoplus_{i=0}^{\infty} \frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2}
\end{equation}
and
\begin{equation}\label{eqn2.10}
{\mathcal O}_{X,p}\cong \bigoplus_{i=0}^{\infty} \frac{m_p^{i}}{m_p^{i+1}}.
\end{equation}
Equation \eqref{eqn2.9} follows from Equation \eqref{eqn2.10}, where the isomorphism of  Equation \eqref{eqn2.10} follows from Taylor
expansion of $f$ in local coordinates $y_1,...,y_n$ at $p$,$f= f_0+f_1+...$, where $f_i\in m_p^i$ is the sum of all the terms of
homogeneous degree $i$.

Under these isomorphisms, we see the  image of $\frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2}$ under the map $$\Omega^1_{X,p} \stackrel{T}{\to} {\mathcal O}_{X,p}$$
is in $\frac{m_p^{i+1}}{m_p^{i+2}}$. Moreover, $$T(\frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2})= \frac{m_p^{i+1}}{m_p^{i+2}}$$  since $m_p^i \cdot m_p=m_p^{i+1}$.

Now from Equation \eqref{eq2.7}, we have the following exact sequence of commutative diagram

\begin{equation}\label{eqn7}
\xymatrix{
0\ar[r]&\ker (T)\ar[r]\ar[d]^{\psi^*} & {\Omega}^1_{X,p}\ar[d]^{\psi^*}\ar[r]^{T}\ar[d]&m_{p}\ar[d]^{\psi^*}\ar[r]&0\\
0\ar[r]&\ker (T_o)\ar[r] & {\Omega}^1_{\C^n,o}\ar[r]^{T_0}& m_{o}\ar[r]&0.
}
\end{equation}

\begin{lemma}\label{lemma8}
There exists

The pullback map ${\psi^*}:{\Omega}^1_{X,p} \to   {\Omega}^1_{\C^n,o}$ is surjective.
\end{lemma}

\begin{proof}
Note that  $\Omega^1_{X,p}\cong \bigoplus_{i=0}^{\infty} \frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2}$
and  $\Omega^1_{\C^n,o}\cong \bigoplus_{i=0}^{\infty} \frac{m_o^i}{m_o^{i+1}}\otimes \frac{m_o}{m_o^2}$.

Since $\dim X=n$, $\psi^*$ induces an isomorphism $\psi^*: T_p^*X\cong T_o^*(\C^n)$ of $\C$-vector spaces, i.e., $\psi^*:\frac{m_p}{m_p^2}\cong  \frac{m_o}{m_o^2}$.
In other words, $\psi^*:\frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2} \to  \frac{m_o^i}{m_o^{i+1}}\otimes \frac{m_o}{m_o^2}$ is an isomorphism for $i=0$.

Now we use inductions on $i$.
Suppose that $\psi^*:  \frac{m_p^i}{m_p^{i+1}}\otimes \frac{m_p}{m_p^2}\to \frac{m_o^i}{m_o^{i+1}}\otimes \frac{m_o}{m_o^2} $ is surjective for $i=k$.

Consider the following commutative diagram
$$\xymatrix{\frac{m_p^k}{m_p^{k+1}}\otimes \frac{m_p}{m_p^2}\ar[r]^-{T_p}\ar[d]^{\psi^*}& \frac{m_p^{k+1}}{m_p^{k+2}}\ar[d]^{\psi^*}\ar[r]&0\\
\frac{m_0^k}{m_0^{k+1}}\otimes\frac{m_o}{m_o^2}\ar[r]^-{T_o}& \frac{m_o^{k+1}}{m_o^{k+2}}\ar[r]&0,
}
$$
we get from Equation \eqref{eqn7} both $T$ and $T_o$ are surjective. By the inductive hypothesis, the first column $\psi^*$ is isomorphism. Hence we
know $\psi^*$ on the second column is surjective. Therefore,
$$\psi^*:\frac{m_p^{k+1}}{m_p^{k+2}}\otimes \frac{m_p}{m_p^2}\to\frac{m_o^{k+1}}{m_o^{k+2}} \otimes \frac{m_o}{m_o^2}$$ is also surjective.
This completes the induction and hence the proof of the lemma.
\end{proof}

\begin{lemma}\label{lemma9}
The pullback map ${\psi^*}:m_p\to m_o$ is an isomorphism.
\end{lemma}
\begin{proof}
Since $\psi^*:m_o\to m_p$ is a prior injective as we explained above, we only need to show it is surjective.
Note that in  Equation \eqref{eqn7}, both
$T$ and $T_o$ are surjective. Moreover, $\psi^*$ on the middle column is surjective by Lemma \ref{lemma8}.
Therefore, $\psi^*$ on the third column is surjective. This completes the proof of the lemma.
\end{proof}

\begin{remark}
Harvey and Lawson applied to their techniques to the flow given by a $\C^*$-action on a compact K\"{a}hler manifold to obtain a complex graph $\mathcal{T}$.
Moreover, they obtained that $\mathcal{T}$ is of finite volume, $\overline{\mathcal{T}}$ are of an analytic subvariety (see \cite{Sommese}).
\end{remark}

\begin{theorem}\label{Thm}
If $X$ is a nonsingular irreducible projective variety admitting a $\C$-action with nonempty isolated fixed point, then $X$
is rational.
\end{theorem}
\begin{proof}
Since $\psi^*:\mathcal{O}_{\C^n,p}\to \mathcal{O}_{X,o}$ is an injective $\C$-algebraic homomorphism by Lemma \ref{lemma2.6}.
By Lemma \ref{lemma9}, $\psi^*: m_p\to m_o$ is  surjective.
Since $\psi^*$ maps constants in $\mathcal{O}_{X,p}$  exactly onto constants in  $\mathcal{O}_{\C^n,o}$ , we obtain that
$\psi^*: \mathcal{O}_{X,p}\to \mathcal{O}_{\C^n,o}$ is also surjective.  Hence $\psi^*:\mathcal{O}_{X,p} \to\mathcal{O}_{\C^n,o}$ is
an isomorphism.
Since the rational function field $K(X)$ (resp. $K(\P^n)$)
of $X$ (resp. $\C^n$) is the fractional field of $\mathcal{O}_{X,p}$ (resp. $\mathcal{O}_{\C^n,o}$),  $\psi^*:K(X)\to K(\C^n)$
is an isomorphism as a field. This implies that $X$ is birational to $\C^n$ (and hence $\P^n$) (see \cite{Hartshorne}) and hence $X$ is rational.
This completes the proof of the theorem.
\end{proof}

\begin{remark}
Since the condition of regularity of $X$ is a key part in the proof of Theorem \ref{Thm} and in the processing to construct equivariant projections, it obviously does not work for varieties
with singularity. Moreover,  cones over nonsingular projective varieties are  counterexamples which prevent  generalizing Theorem \ref{Thm}
to the singular case.
\end{remark}

\section{Uniform Upper bound of Euler numbers and Betti numbers}

In this section, let $X$ be a nonsingular complex projective variety of dimension $n$ admitting a $\C$-action with exactly one fixed point.
Recall that
 \begin{equation}
 I-P=d\circ T+T\circ d:\mathcal{E}^k(X)\to \mathcal{D}'^k(X)
 \end{equation}

For $\alpha \in \mathcal{E}^k(X)$ such that $d \alpha=0$, we have
$$
\alpha-P\alpha= d T\alpha.
$$

Since $P\alpha=\lim_{t\infty}\phi_t^*(\alpha) $, the value of $P\alpha$ at $x\in X$
is given by
\begin{equation}\label{eqn3.2}
(P\alpha)(x)=\lim_{t\to \infty} \phi_t^*(\alpha)(x)=\lim_{t\to \infty}\alpha\circ \phi_t(x)=\alpha(\lim_{t\to \infty}\phi_t(x))=\alpha(p).
\end{equation}

That is to say, $P\alpha$ is a constant value form on $X$.  Note that
the type of $\alpha$ is preserved under the projector operator $P$ from its definition, we get $P\alpha$ is a constant $(p,q)$-form
is $\alpha$ is a closed form of $(p,q)$-type. Recall that Carrell and Lieberman showed that $H^{i,j}(X)=0$ if $i\neq j$ (see \cite{Carrell-Lieberman}), there is no constant $(i,j)$-form on $X$ if $i\neq j$. For $i=j$, a constant $(i,i)$-form is the
$\C$-linear combination of  $dz_{n_1}\wedge...\wedge dz_{n_i}\wedge d\bar{z}_{m_1}\wedge...\wedge d\bar{z}_{m_i}$. The number
of such form is $ (^n_i)\cdot (^n_i)$.  Hence we get the following upper bound of  Hodge numbers $h^{i,j}(X)=\dim H^{i,j}(X)$, Betti numbers $b^i(X)$ and the Euler number $\chi(X)$ of such a
nonsingular complex projective variety $X$.

\begin{theorem}\label{Th3.2}
  Let $X$ be a nonsingular complex projective variety of $\dim X=n$ admitting a $\C$-action with exactly one fixed point.
  Then $b^{2i}(X)= h^{i,i}(X)\leq (^n_i)$ and the Euler number $\chi(X)$ of $X$ is upper bounded by $2^n$.
\end{theorem}
\begin{proof}
  Note that $b^{2i-1}(X)=0$ and $\dim b^{2i}(X)=dim H^{i,i}(X)$, which
   follows from the Hodge decomposition Theorem and Carrell-Lieberman's theorem (\cite{Carrell-Lieberman}) for $X$.
 Moreover,  we have $dim H^{i,i}(X)\leq  (^n_i)$ by Lemma \ref{lemma3.4} below.

  Then we have $$\chi(X)=\sum_{i=0}^n b^{2i}(X)\leq \sum_{i=0}^n (^n_i)=2^n.$$
\end{proof}

\begin{lemma}\label{lemma3.4}
   Let $X$ be a nonsingular complex projective variety of $\dim X=n$ admitting a $\C$-action with exactly one fixed point.
  Then $b^{2i}(X)\leq (^n_i)$.
\end{lemma}
\begin{proof}

Since $b^2(X)=h^{1,1}(X)$, we only need to consider the closed $(1,1)$-form.
 Let  $\alpha$ be a closed $(1,1)$-form.
  Then  we have $\alpha$ is homologous to $P\alpha$, which is
a  $(1,1)$-form of constant value  by Equation \eqref{eqn3.2}. Since a  $(1,1)$-form $P\alpha$
is a $\C$-linear span of $dz_i\wedge d\bar{z}_j$. A real $(1,1)$-form  is
spanned by $dz_i\wedge d\bar{z}_j+dz_j\wedge d\bar{z}_i=
d(z_i+z_j)\wedge d(\bar{z}_i+\bar{z}_j)-dz_i\wedge d\bar{z}_i-dz_j\wedge d\bar{z}_j$
and an imaginary $(1,1)$-form is spanned by
$dz_i\wedge d\bar{z}_j-dz_j\wedge d\bar{z}_i=-\frac{1}{\sqrt{-1}}\{
d(z_i+\sqrt{-1}z_j)\wedge d(\bar{z}_i-\sqrt{-1}\bar{z}_j)-dz_i\wedge d\bar{z}_i-dz_j\wedge d\bar{z}_j\}$.
Hence the class of $\alpha$ is spanned by the form $\xi\wedge \bar{\xi}$, where $\xi$ is a constant $1$-form.
The space spanned constant $1$-forms on $X$  is of dimension at most $n$ since $\dim X=n$.
The method works for $i>2$ to get $b^{2i}(X)\leq(^n_i)$ by induction.
\end{proof}

\begin{remark}
  Lemma \ref{lemma3.4} is a generalization to the ``unexpected consequence" of Akyildiz-Carrell on the upper bound
  of the second Betti number (see \cite[Th2]{Akyildiz-Carrell}). Note that the condition is much weaker than that
  in Akyildiz-Carrell's result.
\end{remark}
\begin{remark}
  This bound is the optimal upper bound for $b^{2i}(-)$ since $X=(\P^1)^n$ admits a $\C$-action with exact one fixed point and
  $b^{2i}((\P^1)^n)=(^n_i)$.
\end{remark}

\begin{remark}
Since a connected nonsingular projective variety $X$ admits $\C$-action with exact one fixed point is equivalent to
that $X$ admits a holomorphic vector field with exact one zero locus by using
a result of Bia{\l}ynicki-Birula \cite{Bialynicki-Birula1},
the condition in Theorem \ref{Th3.2} can be replaced to be  that $X$ is a nonsingular complex projective variety
of $\dim X=n$ admitting a holomorphic vector field with exactly one zero locus.
\end{remark}

\begin{remark}
Let $X$ admit a holomorphic vector field $V$ with exact one zero locus.
  Since the upper of Euler number of $X$  depends only on the $\dim X$, one can apply equivariant blowing ups along zero locus to get nonsingular projective variety $\widetilde{X}$. The Euler number $\chi(\widetilde{X})$  increases in the process but $\dim \widetilde{X}=\dim X$. Hence
  the zero locus of the induced holomorphic vector field on $\widetilde{X}$ will not be isolated. Hence the holomorphic vector field $V$ is never
  ``generic" in the sense of Lieberman \cite{Lieberman1}.
\end{remark}

\begin{proposition}
  Let $Y$ admit a holomorphic vector field $V$ with exact one zero locus. Then $Y$ is rational minimal variety in the sense that $Y$ is not
  the blow up of any smooth projective variety along a codimension at least two subvarieties.
\end{proposition}
\begin{proof}
Assume that $Y=\widetilde{X}$, where
$\sigma:\widetilde{X}\to X$ is  a blow up of a smooth projective variety $X$ along a smooth projective
subvariety of codimension at least two.
Note that a blow up $\sigma:\widetilde{X}\to X$ is always equivariant (see \cite{Lieberman2}).
The exceptional divisor $E$ is $\C$-invariant since its $\widetilde{X}-E\cong X-pt$ is invariant. If there exists a point on
$E$ is not fixed point, then the fixed point set must be of codimension bigger than or equal to 2.

Let  $O_1,O_2$ be two different orbits in $X$ such that the tangent vectors along the orbits are different.
The proper transform  $\widetilde{O}_1,\widetilde{O}_2$ of two different orbits $O_1,O_2$ in $X$ are $\C$-invariant and hence
there is a fixed point on each of  $\widetilde{O}_1,\widetilde{O}_2$. By the property of the blow up construction, we know
$\widetilde{O}_1\cap \widetilde{O}_2=\emptyset$. Then $\widetilde{X}$ have at least two points, contradicts to
the assumption that $Y=\widetilde{X}$ has exact one fixed point.
\end{proof}

\section{Surfaces admits an additive group action  with exact one fixed point}
It is easy to see that $\P^1$ is the only non-singular complex projective curve admitting a holomorphic vector field with
 exact one zero point.
In the section, we give the  classification of complex projective surfaces admitting a holomorphic vector field with exact one zero point.
 We will show that $\P^2$ and $\P^1\times\P^1$ are only non-singular complex projective surfaces with this property.
Let $X$ be a nonsingular complex projective surface.
As we mentioned above, a nonsingular projective surface $X$ admitting a holomorphic vector field with exact one zero point
is equivalent to it admits a $\C$-action with exact one fixed point.

By Theorem \ref{Th3.2}, if $\dim X=2$, then $\chi(X)\leq 4$ and $b^2(X)\leq 2$.
Since a prior $b^2(X)\geq 1$. Topologically, there are at most two types of $X$.
 Both $\P^2$ and $\P^1\times \P^1$ admit a $\C$-action with exact one fixed point.
 They correspond to $b^2(X)=1$ and $b^2(X)=2$.

 \begin{proposition}\label{Prop4.1}
 Let $\phi:\C\times \Sigma_n\to \Sigma_n$ be a $\C$-action over $\Sigma_n$,
 where $\Sigma_n$ is the Hirzebruch surface $\Sigma_n:=\P(\mathcal{O}_{\P^1}\oplus \mathcal{O}_{\P^1}(n))$.
 If $n\geq 1$, then the fixed point set $X^{\C}$  cannot be  exactly one point set.
 \end{proposition}

 \begin{proof}
   Let $\pi:\Sigma_n\to \P^1$ be the natural projection from the Hirzebruch surface $\Sigma_n$ to $\P^1$. Then the fiber
   of $\pi$ is also isomorphic to $\P^1$. Let $\phi_t:=\phi(t,-):X\to X$ be the flow  given by the action. Then
   $\phi_t$ is fiber-preserving(cf. \cite[p.267]{Siu}). Since $n\geq 1$, the zero section $C$ of $\pi$ has negative self-intersection
   $C^2=-n<0$ (cf. \cite[Ch. IV]{Beauville}). Therefore, it is $\C$-invariant. Let $p\in C$ be a fixed point of the
   induced $\C$-action on $C$. Then the fiber $\pi^{-1}(p)$ is $\C$-invariant since the action is fiber-preserving.

   Suppose that $X^{\C}$ contains exact one point, then $\lim_{t\to \infty}\phi_t(\pi^{-1}(q))=p$. Note that $\pi^{-1}(q)$
   is a fiber, $\phi_t(\pi^{-1}(q))$ is also a fiber for $t\in\C$ since $\phi_t$ is fiber-preserving. Hence $\phi_t(\pi^{-1}(p))=\pi^{-1}(p)$ for all $t\in\C$.

   Let $F$ be a fiber such that $p\notin F$. Since $F\cong \P^1$, we denote that isomorphism by $i:\P^1\cong F$.
   Let $\C\times\P^1\to \Sigma_n$ be defined by
   $$
   \Phi:(t, q)\mapsto \psi_t(q).
   $$
This map is an injective morphism since $\psi_t$ is a $\C$-action on $\Sigma_n$.

The map  $\Phi$ is extended to $\P^1\times\P^1$ by $\widetilde{\Phi}(\infty,q):=\lim_{t\to\infty}\psi_t(q)=p$.
Then we get a morphism $\widetilde{\Phi}:\P^1\times\P^1\to \Sigma_n$. This is impossible except for the case
that $n=0$ since there does not exist a birational morphsim from $\P^1\times\P^1$ to a Hirzebruch surface
$\Sigma_n$ for $n>0$. The last statement follows from the fact that their corresponding Betti numbers are the same.
\end{proof}

\begin{example}\label{ex2.2}
  In Proposition \ref{Prop4.1}, $n\geq 1$ is necessary. When $n=0$, then $\Sigma_0=\P^1\times \P^1$.
  Let $\C\times \P^1\times \P^1\to \P^1\times \P^1$ be the $\C$-action given by $(t, [u:v],[x:y])\mapsto ([u+tv:v],[x+ty:y])$.
  One can check that the fixed point  of this action is $([1:0],[1:0])$.
\end{example}

\begin{example}\label{ex2.3}
  It is  well-known  that there exists a $\C$-action $\C\times \P^2\to\P^2$ on $\P^2$  given by $(t,[x:y:z])\mapsto [x+ty+\frac{1}{2}t^2z:y+tz:z]$ whose fixed point is $[1:0:0]$.
\end{example}

Therefore, the only nonsingular complex projective surfaces admitting a $\C$-action with exact one fixed point
  are $\P^2$ and $\P^1\times\P^1$.
In summary, we have the following result classification of nonsingular complex projective surface
 admitting a holomorphic vector field with exact one zero locus.

\begin{theorem}\label{Th4.4}
  Let $X$ be a nonsingular complex projective surface admitting a holomorphic vector field with exact one zero locus. Then
  $X$ is either isomorphic to $\P^2$ or $\P^1\times \P^1$.
\end{theorem}

Clearly, both $\P^2$ and $\P^1\times \P^1$ also admit a $\C^*$-action with isolated fixed points. Hence we also show that
the strong version Carrell conjecture  in dimension two.
\begin{corollary}
   Let $X$ be a nonsingular complex projective surface admitting a holomorphic vector field with exact one zero locus. Then
  $X$  admits a $\C^*$-action with isolated fixed points.
\end{corollary}

\begin{remark}
  In Theorem \ref{Th4.4}, ``nonsingular complex projective surface" can be replaced to be ``compact K\"{a}hler surface"
  since such a compact K\"{a}hler surface is algebraic  (cf. \cite{Carrell-Lieberman}).

\end{remark}
\section{Threefolds admits an additive group action  with exact one fixed point}

  It will be more complicated to classify higher dimensional
  nonsingular complex projective varieties admitting a $\C$-action with exact one fixed point.
  For example, there exist rational Fano threefolds admitting $\C$-action with exact one fixed point but not isomorphic to the product of
  projective spaces. Examples can be found in \cite{Akyildiz-Carrell} and references therein. The classification in three dimensional case
  is closely related to the classification of minimal rational threefolds.

According to Theorem \ref{Th3.2}, we see that $b_2(X)\leq 3$ if $X$ is a threefold admits a $\C$-action with exact one fixed  point.
Equivalently, the Picard number of $X$ is less than or equal to 3.

The following result is about the classification of such an $X$ with ${\rm Pic}(X)=1$.
\begin{theorem}\label{Thm5.1}
   Let $X$ be a nonsingular complex projective threefold admitting a $\C$-action with exact one fixed point. If $Pic(X)=1$, then
  $X$ is isomorphic to one of the following varieties:
  \begin{enumerate}
    \item  $\P^3$;
    \item  A smooth quadric in $\P^4$;
    \item   A section of the Grassmannian $G(2,5)\subset \P^9$ by subspace of codimension $3$;
    \item   A Fano threefold $X$ in $\P^{13}$ with $-K_X^3=22$, $b_3(X)=0$ and $g(X)=12$.
  \end{enumerate}
\end{theorem}
\begin{proof}
  First of all, any variety on the list admits a $\C$-action with exact one fixed point.
The obvious construction  similar to Example  \ref{ex2.3} gives $\C$-action
on $X$ such that $X^{\C}$ is exact one point in Case 1 to Case 3.  The fact that the varieties in
Case 4 to Case 6 admits a $\C$-action with exact one fixed point can be found in \cite{Akyildiz-Carrell}, \cite{Konarski}
  and the references therein.

  Now we will show that only the varieties in the list admits $\C$-action with exact one fixed point.
  According the classification of smooth Fano threefolds of Picard number one, 
 there are  exact four threefolds  with $h^{1,2}(X)=0$ (\cite[p.215]{Parshin-Shafarevich}).
    The following threefolds are in the list:
    \begin{enumerate}
  \item $ \P^3$;
    \item  A smooth quadric in $\P^4$;
    \item   A section of the Grassmannian $G(2,5)\subset \P^9$ by subspace of codimension $3$;
    \item   A Fano threefold $X$ in $\P^{13}$ with $-K_X^3=22$, $b_3(X)=0$ and $g(X)=12$.
  \end{enumerate}
\end{proof}

\begin{remark}
Let $X$ be the threefold admitting a $\C$-action with exact one fixed point. 
We guest that if  ${\rm Pic}(X)=2$, then $X$ is isomorphic to  $\P^2\times \P^1$, while
if ${\rm Pic}(X)=3$, then $X$ is isomorphic to  $\P^1\times \P^1\times\P^1$.
This could be deduced  from the classification of smooth minimal rational threefolds.    
\end{remark}

If a nonsingular projective variety $X$ admits a $\C$-action with exact one fixed point,
then $X$ admits a holomorphic vector field with exact one zero locus.
The inverse is also true. This follows from Lieberman that a holomorphic vector field on $X$
 generating a 1-parameter subgroup of the automorphic group $Aut(X)$ is the product
 of $\C^*$ and at most a copy of  $\C$.  The fixed point set of  a smooth projective variety $X$ with
an action of $\C^*$ has at least two disconnected components. Therefore, Theorem \ref{Thm5.1} can be
restated as follows.

\begin{theorem}\label{Thm5.2}
   Let $X$ be a nonsingular complex projective threefold admitting a holomorphic vector field with exact one zero locus. If ${\rm Pic}(X)=1$, then  $X$ is isomorphic to one of the following varieties:
  \begin{enumerate}
  \item $ \P^3$;
    \item  A smooth quadric in $\P^4$;
    \item   A section of the Grassmannian $G(2,5)\subset \P^9$ by subspace of codimension $3$;
    \item   A Fano threefold $X$ in $\P^{13}$ with $-K_X^3=22$, $b_3(X)=0$ and $g(X)=12$.
  \end{enumerate}
\end{theorem}

\section*{Acknowledgements}
  We would like to thank R. Harvey and B. Lawson for useful comments and clarification  in the early version.
I would thank Yifei Chen for a detailed explanation on the classification of minimal rational threefolds.
This work is partially supported by NSFC(11771305, 11821001).


\end{document}